Bare wire to insulated wire has the ratio 29:41 for a certain grade. The bare wire weighs 87 lbs. Determine weight of insulated.
Help me solve this.

Answer by rfer(16322)

29/87=41/x
29x=87*41
x=3567/29
x=123 lbs